When 50.00 mL of aqueous HCl was mixed with 50.00 mL of NaOH (in large excess), the temperature of the solution increased from 25.00 °C to 32.40 °C. The reaction is
NaOH(aq) + HCl(aq) ↔ NaCl(aq) + H 2O(aq) -- ΔH = -57.3 kJ
What was the molarity of the original HCl solution? Assume the heat capacity of the solution is the same as pure water (4.184 J/g*°C), the density of the solution is 1.00 g/mL and there is no loss of heat to the surroundings.

Key Concept:: Calorimetry and enthalpy change is an extensive property. V, d → m solution; ΔT, heat capacity → heat; ΔH→ mol of HCl; V → [HCl]
1
Expert's answer
2020-12-15T07:50:00-0500

∆H = nC∆T

-57300 = n * 4.18 *36.5 * (32.40-25)

-57300 = n * 4.18*36.5*7.40


57300 = n× 1129.01

n = 57300/1129.01

n = 50 moles

molarity = no. of moles/volume

molarity = 50moles/0.05L = 1000M